Lucy is 4 times as old as Amy. Three years ago, Lucy was 13 times as old as Amy. How old is Lucy now?
Answer by checkley77(12844)

L=4A
4A-3=13(A-3)
4A-3=13A-39
4A-13A=-39+3
-9A=-36
A=-36/-9
A=4 IS THE AGE OF AMY.
L=4*4=16 IS THE AGE OF LUCY.
PROOF:
16-3=13(4-3)
13=13*1
13=13